A simple random sample of size n = 12 is obtained from a population with u = 69 and o = 14. (a) What must be true regarding tB and C
Answer #1

Solution :

Given that,

mean = \mu = 69

standard deviation = \sigma = 14

n = 12

The sampling distribution of mean is ,

\mu\bar x = \mu = 69

The sampling distribution of standard deviation is ,  

\sigma\bar x = \sigma / \sqrt n = 14 / \sqrt 12 = 4.0415

b)

P(x \leq 72.3)

= P[(x - \mu ) / \sigma \leq (72.3 - 69) /4.0415 ]

= P(z \leq 0.82)

= 0.7939

probability = 0.7939

c)

P(x \geq 70.2 ) = 1 - P(x  \leq 70.2)

= 1 - P[(x - \mu ) / \sigma \leq (70.2 - 69) /4.0412 ]

= 1 -  P(z \leq 0.30 )  

= 1 - 0.6179 = 0.3821

probability = 0.3821